The population for a clinical study has 500 Asian, 1000 Hispanic and 500 Native American people. What is good way of sampling this population to ensure that the distribution of various sub- populations is maintained if only 100 samples have to be chosen? Give the distribution of the various sub-populations in the final sample.
Answer #1

CONCEPT

Proportionate sampling is a sampling strategy (a method for gathering participants for a study) used when the population is composed of several subgroups that are vastly different in number. The number of participants from each subgroup is determined by their number relative to the entire population.

No of samples from a particular group= Population of that particular group/ total population

Answer- Good way of sampling this population to ensure that the distribution of various sub- populations is maintained is STRATIFIED SAMPLING (As the population is divided into different strata i.e, Asian, Hispanic and native Americans)

   TOTAL POPULATION= 500+ 1000+ 500= 2000

  IF ONLY !00 SAMPLES HAVE TO BE CHOOSEN, THEN WE WILL USE PROPORTIONATE SAMPLING TO DECIDE HOW MaNY SAMPLES ARE TO BE DRAWN FROM EACH STRATA

Number of Asian in the sample= (500/2000) * 100 =25

Number of Hispanic in the sample= (1000/2000)* 100= 50

Number of Native American in the sample= (500/2000) * 100= 25

THUS we will have 25 Asians, 50 Hispanic and 25 Native Americans in the final sample which will be selected randomly from each strata